# Errors-In-Variables Model Fitting for Partially Unpaired Data Utilizing Mixture Models

### v1

@@ Appendix F Application of $R_{\delta}^{2}$
In [20] a consistent goodness of fit measure for errors-in-variables multiple linear regression is presented if the standard deviations of the input variables are known which direct application is presented in this appendix. With $X$ the $L\times k$-Matrix of $L$ observations and $k$ predictor variables, $\boldsymbol{y}$ the $L\times 1$ vector of output observations, $\Sigma_{\delta}$ the $k\times k$ covariance matrix of predictor variables, $S=\frac{1}{L}\,X^{T}\,P\,X$ and $P=I_{L\times L}-\frac{1}{L}\mathbf{1}_{L\times L}$ ($\mathbf{1}$ being the matrix consisting of $1$s) the goodness of fit is defined by
@@
where $\boldsymbol{b}$ is the vector of fitted slopes, i.e. in the notation of the multiple linear regression model of this paper $\boldsymbol{b}=(\alpha_{2},\alpha_{3},\dots,\alpha_{k})^{T}$ being independent of the intercept $\alpha_{1}$.

### v2

@@ Appendix D Application of $R_{\delta}^{2}$
In [Cheng et al., 2014] a consistent goodness of fit measure for errors-in-variables multiple linear regression is presented if the standard deviations of the input variables are known. We show its direct application in this appendix. With $X$ the $L\times k$-Matrix of $L$ observations and $k$ predictor variables, $\boldsymbol{y}$ the $L\times 1$ vector of output observations, $\Sigma_{\delta}$ the $k\times k$ covariance matrix of predictor variables, $S=\frac{1}{L}\,X^{T}\,P\,X$ and $P=I_{L\times L}-\frac{1}{L}\mathbf{1}_{L\times L}$ ($\mathbf{1}$ being the matrix consisting of $1$s) the goodness of fit is defined by
@@
where $\boldsymbol{b}$ is the vector of fitted slopes, i.e., in the notation of the multiple linear regression model of this paper $\boldsymbol{b}=(\alpha_{2},\alpha_{3},\dots,\alpha_{k})^{T}$ being independent of the intercept $\alpha_{1}$.